MEASUREMENT OF EXCITED STATE DIPOLE MOMENTS OF 7-AZAINDOLE IN THE GAS PHASE.^{a}

\underline{CHEOLHWA KANG}, JOHN T. YI and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

The rotationally resolved S_{1} ← S_{0} fluorescence excitation spectrum of 7-azaindole has been obtained in the presence of an external electric field. The Stark effect on the rotational energy levels allows us to accurately determine the values for each component of the dipole moment in both electronic states. The magnitudes of dipole moment components are μ_{a} (S_{0}) = 1.45 ± 0.01 D, μ_{b} (S_{0}) = 0.65 ± 0.06 D and μ_{a} (S_{1}) = 2.23 ± 0.01 D, μ_{b} (S_{1}) = 0.55 ± 0.07 D. The large increase in μ_{a} is a result of changes in the electron distribution in the S_{1} state. A detailed discussion of these changes will be given.

aWork supported by NSF.